OpenStudy (anonymous):

12 question review need help fan and medal :)

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Hold up give me a sec

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

A favor: Post just one question at a time, please. Thank you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The ordered pairs (1,3), (2,9), (3,27), (4,81) and (5,243) represent a function. What is a rule that represents this function? Y=x^3 Y=3x y=3^x y=x+2

OpenStudy (anonymous):

ar u in 6th

OpenStudy (anonymous):

no lol just didn't study this stuff was in the hospital with a broken neck

OpenStudy (anonymous):

y = (x+4)^2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Wait how do I put that togethor?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Wait i'm sorry confused already :D

OpenStudy (anonymous):

ooh sorry

OpenStudy (anonymous):

So, Do you think it's D I'm kind of not getting it

OpenStudy (anonymous):

no its c

OpenStudy (anonymous):

OH thanks, Next question?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The ordered pairs (1,25),(2,36),(3,49),(4,64) and (5,81) represent a function. What is a rule that represents this function? Y=(5x)^2 Y=x+11 y+x^5 y=(x+4)^2

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

hmm.. Is this hard for you to?

OpenStudy (anonymous):

what garde are you in

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

10th grade

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

i looked this one up and its d

OpenStudy (anonymous):

im in 7th grade but really good at math

OpenStudy (anonymous):

How did you look it up?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Suppose yogurt is on sale for $0.50 a piece, and you have a coupon for $0.30 off your total purchase. Write a function rule for the cost of n yogurts. How much would 20 yogurts cost? C(n)=0.5-0.3;$9.70 C(n)=0.3n-0.5;$5.50 C(n)=0.2n;$4.00 C(n)=0.5n;$10.00

OpenStudy (anonymous):

cn=0.5-0.3;$9.70

OpenStudy (anonymous):

or c20

OpenStudy (anonymous):

Ok sorry I was trying to take a picture of the next question

OpenStudy (anonymous):

it ok

OpenStudy (anonymous):

https://gyazo.com/df520971128ec898e55fd36065ff02bd

OpenStudy (anonymous):

idk im sorry

OpenStudy (anonymous):

these ones are confusing to me

OpenStudy (anonymous):

Well are you up for another maybe?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i hate mapping diagrams

OpenStudy (anonymous):

The function b(n)=6n represents the number of light bulbs b(n) that are needed for n chandeliers. How many lightbulbs are needed for 15 chandeliers 90 lightbulbs 2 lightbulbs 96 lightbulbs 80 lightbulbs

OpenStudy (anonymous):

Haha me to i hate mapping diagrams

OpenStudy (anonymous):

80 light bulbs

OpenStudy (anonymous):

thanks :) next lol

OpenStudy (anonymous):

LIsa earns $7.50 per hour mowing lawns a. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. b. How much money does lisa earn if she works for 5 hours and 45 minutes? m(h)=5h+45;$37.95 m(h)=7.50h;$40.88 m(h)=h/7.50;$1.30 m(h)=7.50h;$43.13

OpenStudy (anonymous):

My wrist hurts now from typing that out lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this is a snap m(h)=7.5(h)

OpenStudy (anonymous):

So, I think you mean m(h)=7.50h;43.13

OpenStudy (anonymous):

A magician charges $50.00 for a visit and an additional $7.50 for each hour he performs. The function rule C=7.50h+50.00 describes the relationship between the number of hours h and the total cost of the visit C. If the magician will only visit a maxiumum of 8 hours, what is a reasonable graph of the function rule?

OpenStudy (anonymous):

yes im lazy :D

OpenStudy (anonymous):

https://gyazo.com/338f22c867f3a2b8f382f37c4cfec021

OpenStudy (anonymous):

Another graph :D

OpenStudy (anonymous):

its the third one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

just a few more hang in there :D

OpenStudy (anonymous):

I have been going through some of them to answer them myself so it makes it more easier on you :)

OpenStudy (anonymous):

This question seems very hard and complicating

OpenStudy (anonymous):

spill it

OpenStudy (anonymous):

:D

OpenStudy (mathmale):

Jerry: Again I ask you to post only one question at a time. Please just close this discussion and type a new question into the "Ask a question" box.

OpenStudy (anonymous):

We're almost done

OpenStudy (anonymous):

It's just him and me working togethor please let us finish

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

https://gyazo.com/9ad5eb31f3dd000dd9180e75cfd50f66

OpenStudy (anonymous):

btw are you in k12

OpenStudy (anonymous):

Are you in k12?

OpenStudy (anonymous):

I'm not

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just asking

OpenStudy (anonymous):

Oh hey in k12 do they have portfolios?

OpenStudy (anonymous):

Or like where you have to do alot of projects?

OpenStudy (anonymous):

yes and no

OpenStudy (anonymous):

I'm in connections academy :) But I was thinking of going to k12 next year

OpenStudy (anonymous):

Yes meaning you do have portfolios?

OpenStudy (anonymous):

i have offline pages and online and offline tests

OpenStudy (anonymous):

this question is tuff

OpenStudy (anonymous):

What? you mean you have offline tests

OpenStudy (anonymous):

yes but some are fun

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

like art and they do microscope assinments

OpenStudy (anonymous):

last year they sent me a rock thing with minerals like sulfer

OpenStudy (anonymous):

Alright lol back to the question now we just went off topic

OpenStudy (anonymous):

4 more questions!!! And i'm done yeah

OpenStudy (anonymous):

Oh really?

OpenStudy (anonymous):

Now I want to go to k12

OpenStudy (anonymous):

idk about that question its tuff

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

Well try because I have to explain it and I have no idea on how to do it

OpenStudy (anonymous):

are you on a computer

OpenStudy (anonymous):

More then one value y for a value of x. Therefore it is a function

OpenStudy (anonymous):

Yeah i'm on a computer

OpenStudy (anonymous):

Is that good for that answer?

OpenStudy (anonymous):

then copy the question and look it up